# MagicQuill: An Intelligent Interactive Image Editing System

### v1

@@ Abstract
As a highly practical application, image editing encounters a variety of user demands and thus prioritizes excellent ease of use. In this paper, we unveil MagicQuill, an integrated image editing system designed to support users in swiftly actualizing their creativity. Our system starts with a streamlined yet functionally robust interface, enabling users to articulate their ideas (e.g., inserting elements, erasing objects, altering color, etc.) with just a few strokes. These interactions are then monitored by a multimodal large language model (MLLM) to anticipate user intentions in real time, bypassing the need for prompt entry. Finally, we apply the powerful diffusion prior, enhanced by a carefully learned two-branch plug-in module, to process the editing request with precise control. Please visit https://magic-quill.github.io to try out our system.
@@ 2.1 Image Editing
Image editing involves modifying the visual appearance, structure, or elements of an existing image [19]. Recent breakthroughs in diffusion models [17, 49, 44] have significantly advanced visual generation tasks, outperforming GAN-based models [15] in terms of image editing capabilities. To enable control and guidance in image editing, a variety of approaches have emerged, leveraging different modalities such as textual instructions [47, 14, 6, 65, 32, 11], masks [20, 48, 69, 23], layouts [68, 10, 33], segmentation maps [62, 35], and point-dragging interfaces [37, 38, 36]. Despite these advances, these methods often fall short when precise modifications at the regional level are required, such as alterations to object shape, color, and other details. Among the various methods, sketch-based editing approaches [64, 34, 61, 59, 42, 22, 25] offer users a more intuitive and precise means of interaction. However, the current methods remain limited by the accuracy of the text signals input alongside the sketches, making it challenging to precisely control the information of the editing areas, such as color. To achieve precise control, we introduce two types of local guidance based on brushstrokes: scribble and color, thereby enabling fine-grained control over shape and color at the regional level.
@@ 2.2 MLLMs for Image Editing
Multi-modal large language models (MLLMs) extend LLMs to process both text and image content [16], enabling text-to-image generation [52, 53, 58, 28, 9], prompt-refinement [60, 63], and image quality evaluation [51].
@@ 3.1 Editing Processor
Control Condition from Brushstroke Signal. Let $\mathbf{M}_{add}$ and $\mathbf{M}_{sub}$ denote the binary masks corresponding to add and subtract brush respectively. These masks share the same dimensions as the original image $\mathbf{I}$, where values are set to $1$ in regions corresponding to user brush strokes and 00 elsewhere. The subtract brush masks out the edges from the edge map $\mathbf{E}$, which is initially extracted from the original image using a pre-trained CNN $f_{CNN}$. Conversely, the add brush introduces new edges by setting designated regions to white in the edge map. The resulting modified edge map $\mathbf{E}_{cond}$ serves as the control condition for manipulating geometric structure in the editing processor. This can be formally expressed as
@@
Controllable Image Inpainting. The inpainting branch adopts the UNet [23, 45] architecture, incorporating the masked image feature into the pre-trained diffusion network. This branch inputs the concatenated noisy latent at $t$-th step $z_{t}$, masked image latent $z_{masked}$ extracted using VAE [26] from $\mathbf{I}_{masked}$, and downsampled mask $\mathbf{m}$ by cubic interpolation from $\mathbf{M}$. The inpainting branch processes these features, utilizing a trainable clone of the diffusion model, stripped of cross-attention layers to focus solely on the image feature. The extracted features carrying pixel-level information are inserted into each layer of the frozen diffusion model through zero-convolution layers $\mathcal{Z}$ [66]. Given text condition $\tau$, timestep $t$, let $F(z_{t},t,\mathbf{c};\Theta)_{i}$ represents the feature of the $i$-th layer in the total $n$ layers of the diffusion UNet with parameter $\Theta$. Similarly, let $F^{I}([z_{t},z_{masked},\mathbf{m}],t;\Theta^{I})_{i}$ represents the output of the $i$-th layer in the inpainting UNet, where []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation operation. This feature insertion can be represented by
@@ 3.2 Painting Assistor
Prompt formatting. In our system, we implement two types of question answering (Q&A) [3] tasks to facilitate the Draw&Guess. For the add brush, we utilize a prompt structured as follows: “This is a ‘draw and guess’ game. I will upload an image containing some strokes. To help you locate the strokes, I will give you the normalized bounding box coordinates of the stokes where their original coordinates are divided by the padded image width and height. The top-left corner of the bounding box is at $({x_{1}},{y_{1}})$, and the bottom-right corner is at $({x_{2}},{y_{2}})$. Now tell me in a single word a phrase, what am I trying to draw with these strokes in the image?” The Q&A output directly serves as the predicted prompt. For the subtract brush, we bypass the Q&A process, as the results demonstrate that prompt-free generation achieves satisfactory results.
@@
Dataset Construction. We selected the Densely Captioned Images (DCI) dataset [54] as our primary source. Each image within the DCI dataset has detailed, multi-granular masks, accompanied by open-vocabulary labels and rich descriptions. This rich annotation structure enables the capture of diverse visual features and semantic contexts.
@@
MLLM Fine-Tuning. Our dataset construction method effectively prepares the model to understand and predict user edits, which contains a total of $24,315$ images, categorized under $4,412$ different labels, ensuring a broad spectrum of data for training. To optimize the performance of the MLLM over Draw&Guess, we fine-tuned the LLaVA model, leveraging the Low-Rank Adaptation (LoRA) [18] technique, allowing the efficient fine-tuning without extensively large dataset. Consistent with the original LLaVA training objectives, our approach aims to maximize the likelihood of the correct labels given the input corpora $u$, which is defined as
@@ 3.3 Idea Collector
Interface Design. The user interface of MagicQuill is designed for an intuitive and streamlined image editing experience, as depicted in Figure 2. The interface is divided into several interactive sections, emphasizing ease of use while providing flexible control over the editing process. The interface comprises several key areas: a Prompt Area (A) displaying MLLM-suggested prompts, a Toolbar (B) with essential editing tools, Layer Management (C) for organizing brush strokes, the main Canvas (D) for editing, a Generated Images area (E) for previewing results, Execute Button (F), and Parameter Adjustment (G).
Cross-Platform Support. We implement the Idea Collector as a modular ReactJS component library, designed for cross-platform compatibility with various generative AI frameworks, such as Gradio [1] and ComfyUI [7]. The architecture separates client-side user interactions from server-side model computations through HTTP protocols, enabling platform-independent deployment via standard HTML rendering.
@@ 4.1 Controllable Generation
To thoroughly evaluate the controllable generation capabilities of our editing processor, we compared it with four representative baselines from different categories: (1) SmartEdit [20], an instruction-based editing method. We utilize LLaVA-Next [30] to generate the editing instruction; (2) SketchEdit [64], a GAN-based sketch-conditioned method; (3) BrushNet [23], the mask and prompt-guided inpainting method; and (4) a composite baseline combining BrushNet [23] and ControNet [66]. As illustrated in Fig. 6, the instruction-based method, SmartEdit, tends to produce outputs that are too random, lacking the precision required for accurate editing purposes. Similarly, while BrushNet enables region-specific modifications, it struggles with maintaining predictable detail generation even with ControlNet enhancement, making precise manipulation challenging. In contrast, our model achieves more accurate edge alignment and color fidelity, which we attribute to our specialized design of the inpainting and control branch that emphasizes these aspects.
@@
We further conducted a quantitative analysis of our constructed test dataset in Sec. 3.2, which contains $490$ images. Our model outperformed the baselines across all key metrics as in Tab. 1. These results demonstrate significant improvements in controllable generation.
4.2 Prediction Accuracy
@@
To further qualitatively evaluate the Painting Assistor, we conducted a user study with $30$ participants who freely edited images using our system. Participants rated the Painting Assistor on a $5$-point scale for prediction accuracy ($1$: very poor, $5$: excellent) and efficiency facilitation ($1$: significantly reduced, $5$: significantly enhanced). As shown in Fig. 7, $86.67\%$ of users rated prediction accuracy at least $4$, validating the ability of our fine-tuned MLLM to interpret user intentions. Similarly, $90\%$ rated efficiency facilitation $4$ or above, confirming that Draw&Guess effectively streamlines the editing process by reducing manual prompt inputs. The average scores for accuracy and efficiency were $4.07$ and $4.37$.
@@ 4.3 Idea Collection Effectiveness and Efficiency
Baseline. The baseline system was implemented as a customized ComfyUI workflow, replacing our Idea Collector interface with an open-source canvas, Painter Node [40]. This setup enables the focus on the value provided with our Idea Collector by controlling other variables.
Procedure. The study lasted approximately $30$ minutes for each participant with two systems (our system and the baseline). Each session began with a brief introduction to the system using the case illustrated in Fig. 1. Participants then had $5$ minutes to freely explore and edit images. After using both systems, participants completed a questionnaire with $22$ questions ($10$ questions per system covering all four dimensions and $2$ questions regarding the Painting Assistor detailed in Sec. 4.2). We employed the System Usability Scale (SUS) [5] for scoring, using a Likert scale from $1$ (strongly disagree) to $5$ (strongly agree), to capture a global view of subjective usability for each system.
Figure 8: Comparative user ratings between our system and the baseline in four dimensions, with standard deviation shown as error bars.
@@ 5 Conclusion
In conclusion, our interactive image editing system MagicQuill effectively addresses the challenges of performing precise and efficient edits by combining the strengths of the Editing Processor, Painting Assistor, and Idea Collector. Our comprehensive evaluations demonstrate significant improvements over existing methods in terms of controllable generation quality, editing intent prediction accuracy, and user interface efficiency. For future work, we aim to expand the capabilities of our system by incorporating additional editing types, such as reference-based editing, which would allow users to guide modifications using external images. We also plan to implement layered image generation to provide better editing flexibility and support for complex compositions. Moreover, enhancing typography support will enable more robust manipulation of textual elements within images. These developments will further enrich our framework, offering users a more versatile and powerful tool for creative expression in digital image editing.
@@
Appendix A Implementation Details
A.1 Editing Processor
Our Editing Processor is built upon Stable Diffusion v1.5 [44] and is compatible with all customized fine-tuned weights. We set the control parameters with inpainting strength $w_{I}=1.0$ and control strength $w_{C}=0.5$, while expanding the mask region by $15$ pixels during controllable inpainting. The generation process employs the Euler ancestral sampler with Karras scheduler [24], requiring $20$ steps per generation. On standard hardware, generating a $512\times 512$ resolution image takes approximately $2$ seconds with $15$ GB VRAM consumption. For the control branch, we conduct fine-tuning on the LAION-Aesthetics dataset [46], specifically selecting images with aesthetic scores above $6.5$. The training process spans $3$ epochs with a learning rate of $65e-$ 6 and batch size of $8$.
A.2 Painting Assistor
We fine-tune a LLaVA-1.5 model with $7B$ parameters for Draw&Guess task on our own constructed dataset in Sec. 3.2, leveraging LoRA [18]. The LoRA rank and alpha are $64$ and $16$ respectively. The model is trained for $3$ epochs with a learning rate of $52e-$ 5 and batch size of $8$. Under $4$-bit quantization, the model achieves real-time prompt inference within $0.3$ seconds using only $5$ GB VRAM, enabling efficient on-the-fly prompt generation with satisfactory accuracy.
A.3 Idea Collector
Cross-platform Support. Besides Gradio, MagicQuill can also be integrated into ComfyUI as a custom node, as shown in Fig. 9. It is designed with customizable widgets for parameter settings and extensible architecture for future platform integrations.
@@ A.3 Idea Collector
Usage Scenario. To demonstrate the user-friendly workflow of MagicQuill, we present an illustrative scenario: A user wants to modify an image of a complete cake, cutting a slice out of it, as shown in Fig 2. The user begins by uploading the image through the toolbar, which provides access to a range of tools (Fig. 2-B). Using the add brush, the user outlines the slice to be cut directly on the canvas (Fig. 2-D). Meanwhile, the Draw & Guess feature introduced in Sec. 3.2 predicts that the user intends to manipulate a “cake” and suggests the relevant prompt automatically in the prompt area (Fig. 2-A). Afterward, the user switches to the subtract brush to fill in the outlined slice, visually marking the area to be removed from the cake. For additional precision, the eraser tool is available to refine the cut. Once the adjustments are made, the user generates the image by clicking the Run button (Fig. 2-F), which runs the model detailed in Sec. 3.1.
@@
Appendix B Failure Case
B.1 Failure Case of Editing Processor
Scribble-Prompt Trade-Off. We observe quality degradation when user-provided add brush strokes deviate from the semantic content specified in the prompt, a common occurrence among users with limited artistic skills. This creates a fundamental trade-off: strictly following the scribble structure may compromise the generation quality with respect to the text prompt. To address this issue, we propose adjusting the edge control strength.

### v2

@@ Abstract
As a highly practical application, image editing encounters a variety of user demands and thus prioritizes excellent ease of use. In this paper, we unveil MagicQuill, an integrated image editing system designed to support users in swiftly actualizing their creativity. Our system starts with a streamlined yet functionally robust interface, enabling users to articulate their ideas (e.g., inserting elements, erasing objects, altering color, etc.) with just a few strokes. These interactions are then monitored by a multimodal large language model (MLLM) to anticipate user intentions in real time, bypassing the need for prompt entry. Finally, we apply the powerful diffusion prior, enhanced by a carefully learned two-branch plug-in module, to process the editing request with precise control. Please visit the project page to try out our system.
@@ 2.1 Image Editing
Image editing involves modifying the visual appearance, structure, or elements of an existing image [24]. Recent breakthroughs in diffusion models [22, 58, 53] have significantly advanced visual generation tasks, outperforming GAN-based models [20] in terms of image editing capabilities. To enable control and guidance in image editing, a variety of approaches have emerged, leveraging different modalities such as textual instructions [56, 19, 8, 80, 15, 72, 18, 5, 73, 29, 31, 10, 4], masks [25, 57, 84, 28, 67], layouts [83, 14, 40], segmentation maps [77, 43], strokes [76, 44], references [39, 41, 59, 9], and point-dragging interfaces [46, 47, 45]. Despite these advances, these methods often fall short when precise modifications at the regional level are required, such as alterations to object shape, color, and other details. Among the various methods, sketch-based editing approaches [79, 42, 75, 71, 27, 32, 51] offer users a more intuitive and precise means of interaction. However, the current methods remain limited by the accuracy of the text signals input alongside the sketches, making it challenging to precisely control the information of the editing areas, such as color. To achieve precise control, we introduce two types of local guidance based on brushstrokes: scribble and color, thereby enabling fine-grained control over shape and color at the regional level.
@@ 2.2 MLLMs for Image Editing
Multi-modal large language models (MLLMs) extend LLMs to process both text and image content [21], enabling text-to-image generation [63, 64, 35, 13], prompt-refinement [74, 78], and image quality evaluation [62].
@@ 3.1 Editing Processor
Control Condition from Brushstroke Signal: Let $\mathbf{M}_{add}$ and $\mathbf{M}_{sub}$ denote the binary masks corresponding to add and subtract brush respectively. These masks share the same dimensions as the original image $\mathbf{I}$, where values are set to $1$ in regions corresponding to user brush strokes and 00 elsewhere. The subtract brush masks out the edges from the edge map $\mathbf{E}$, which is initially extracted from the original image using a pre-trained CNN $f_{CNN}$. Conversely, the add brush introduces new edges by setting designated regions to white in the edge map. The resulting modified edge map $\mathbf{E}_{cond}$ serves as the control condition for manipulating geometric structure in the editing processor. This can be formally expressed as
@@
Controllable Image Inpainting: The inpainting branch adopts the UNet [28, 54] architecture, incorporating the masked image feature into the pre-trained diffusion network. This branch inputs the concatenated noisy latent at $t$-th step $z_{t}$, masked image latent $z_{masked}$ extracted using VAE [33] from $\mathbf{I}_{masked}$, and downsampled mask $\mathbf{m}$ by cubic interpolation from $\mathbf{M}$. The inpainting branch processes these features, utilizing a trainable clone of the diffusion model, stripped of cross-attention layers to focus solely on the image feature. The extracted features carrying pixel-level information are inserted into each layer of the frozen diffusion model through zero-convolution layers $\mathcal{Z}$ [81]. Given text condition $\tau$, timestep $t$, let $F(z_{t},\tau,t;\Theta)_{i}$ represents the feature of the $i$-th layer in the total $n$ layers of the diffusion UNet with parameter $\Theta$. Similarly, let $F^{I}([z_{t},z_{masked},\mathbf{m}],t;\Theta^{I})_{i}$ represents the output of the $i$-th layer in the inpainting UNet, where []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation operation. This feature insertion can be represented by
@@ 3.2 Painting Assistor
Prompt formatting: In our system, we implement two types of question answering (Q&A) [3] tasks to facilitate the Draw&Guess. For the add brush, we utilize a prompt structured as follows: “This is a ‘draw and guess’ game. I will upload an image containing some strokes. To help you locate the strokes, I will give you the normalized bounding box coordinates of the stokes where their original coordinates are divided by the padded image width and height. The top-left corner of the bounding box is at $({x_{1}},{y_{1}})$, and the bottom-right corner is at $({x_{2}},{y_{2}})$. Now tell me in a single word a phrase, what am I trying to draw with these strokes in the image?” The Q&A output directly serves as the predicted prompt. For the subtract brush, we bypass the Q&A process, as the results demonstrate that prompt-free generation achieves satisfactory results.
@@
Dataset Construction: We selected the Densely Captioned Images (DCI) dataset [65] as our primary source. Each image within the DCI dataset has detailed, multi-granular masks, accompanied by open-vocabulary labels and rich descriptions. This rich annotation structure enables the capture of diverse visual features and semantic contexts.
@@
MLLM Fine-Tuning: Our dataset construction method effectively prepares the model to understand and predict user edits, which contains a total of $24,315$ images, categorized under $4,412$ different labels, ensuring a broad spectrum of data for training. To optimize the performance of the MLLM over Draw&Guess, we fine-tuned the LLaVA model, leveraging the Low-Rank Adaptation (LoRA) [23] technique, allowing the efficient fine-tuning without extensively large dataset. Consistent with the original LLaVA training objectives, our approach aims to maximize the likelihood of the correct labels given the input corpora $u$, which is defined as
@@ 3.3 Idea Collector
The user interface of MagicQuill is designed for an intuitive and streamlined image editing experience, as depicted in Figure 2. The interface is divided into several interactive sections, emphasizing ease of use while providing flexible control over the editing process. The interface comprises several key areas: a Prompt Area (A) displaying MLLM-suggested prompts, a Toolbar (B) with essential editing tools, Layer Management (C) for organizing brush strokes, the main Canvas (D) for editing, a Generated Images area (E) for previewing results, Execute Button (F), and Parameter Adjustment (G).
@@ 4.1 Controllable Generation
To thoroughly evaluate the controllable generation capabilities of our editing processor, we compared it with four representative baselines from different categories: (1) SmartEdit [25], an instruction-based editing method. We utilize LLaVA-Next [37] to generate the editing instruction; (2) SketchEdit [79], a GAN-based sketch-conditioned method; (3) BrushNet [28], the mask and prompt-guided inpainting method; and (4) a composite baseline combining BrushNet [28] and ControlNet [81]. As illustrated in Fig. 6, the instruction-based method, SmartEdit, tends to produce outputs that are too random, lacking the precision required for accurate editing purposes. Similarly, while BrushNet enables region-specific modifications, it struggles with maintaining predictable detail generation even with ControlNet enhancement, making precise manipulation challenging. In contrast, our model achieves more accurate edge alignment and color fidelity, which we attribute to our specialized design of the inpainting and control branch that emphasizes these aspects.
@@
We conducted a quantitative analysis of our constructed test dataset in Sec. 3.2, which contains $490$ images. Our model outperformed the baselines across all key metrics as in Tab. 1. These results demonstrate significant improvements in controllable generation.
We additionally compared two stroke-based editing methods SDEdit [44] and UniPaint [76], and the qualitative results are shown below in Fig. 7.
Figure 7: Visual comparison with stroke-based editing baselines.
4.2 Prediction Accuracy & Efficiency Facilitation
@@
To qualitatively evaluate the Painting Assistor, we conducted a user study with $30$ participants who freely edited images using our system. Participants rated the Painting Assistor on a $5$-point scale for prediction accuracy ($1$: very poor, $5$: excellent) and efficiency facilitation ($1$: significantly reduced, $5$: significantly enhanced). As shown in Fig. 8, $86.67\%$ of users rated prediction accuracy at least $4$, validating the ability of our fine-tuned MLLM to interpret user intentions. Similarly, $90\%$ rated efficiency facilitation $4$ or above, confirming that Draw&Guess effectively streamlines the editing process by reducing manual prompt inputs. The average scores for accuracy and efficiency were $4.07$ and $4.37$. We further provide a quantitative analysis with 10 users performing 10 edits, showing an average time savings of 24.92% on iPad per edit and 19.58% on PC per edit, as in the Tab. 4.2.
[…]
w. Paint. Assit. w.o. Paint. Assit. w. Paint. Assit. w.o. Paint. Assit.
13.29s 17.70s (+4.41s) 12.49s 15.53s (+3.04s)
Table 3: Editing Time Comparison w./w.o. Painting Assistor.
@@ 4.3 Idea Collection Effectiveness and Efficiency
Baseline: The baseline system was implemented as a customized ComfyUI workflow, replacing our Idea Collector interface with an open-source canvas, Painter Node [49]. This setup enables the focus on the value provided with our Idea Collector by controlling other variables.
Procedure: The study lasted approximately $30$ minutes for each participant with two systems (our system and the baseline). Each session began with a brief introduction to the system using the case illustrated in Fig. 1. Participants then had $5$ minutes to freely explore and edit images. After using both systems, participants completed a questionnaire with $22$ questions ($10$ questions per system covering all four dimensions and $2$ questions regarding the Painting Assistor detailed in Sec. 4.2). We employed the System Usability Scale (SUS) [7] for scoring, using a Likert scale from $1$ (strongly disagree) to $5$ (strongly agree), to capture a global view of subjective usability for each system.
Figure 9: Comparative user ratings between our system and the baseline, with standard deviation shown as error bars.
@@ 5 Conclusion
In conclusion, our interactive image editing system MagicQuill effectively addresses the challenges of performing precise and efficient edits by combining the strengths of the Editing Processor, Painting Assistor, and Idea Collector. Our comprehensive evaluations demonstrate significant improvements over existing methods in terms of controllable generation quality, editing intent prediction accuracy, and user interface efficiency. For future work, we aim to expand the capabilities of our system by incorporating additional editing types, such as reference-based editing, which would allow users to guide modifications using external images. We also plan to implement layered image generation to provide better editing flexibility and support for complex compositions. Moreover, enhancing typography support will enable more robust manipulation of textual elements within images. These developments will further enrich our framework, offering users a more versatile and powerful tool for creative expression. The system is available at https://magic-quill.github.io.
Acknowledgments. This work was supported by the Research Grant Council of the Hong Kong Special Administrative Region under grant number 16212623 and the Ant Group Research Intern Program.
@@
6 Implementation Details
6.1 Editing Processor
Our Editing Processor is built upon Stable Diffusion v1.5 [53] and is compatible with all customized fine-tuned weights. We set the control parameters with inpainting strength $w_{I}=1.0$ and control strength $w_{C}=0.5$ for both edge and color control signals, while expanding the mask region by $15$ pixels during controllable inpainting. We use separate ControlNets to independently control edge and color. Although the two signals may conflict, our model blends them using adjustable control weights ($0.5$ by default), allowing users to achieve more precise control. The generation process employs the Euler ancestral sampler with Karras scheduler [30], requiring $20$ steps per generation. On standard hardware, generating a $512\times 512$ resolution image takes approximately $3$ seconds with $10$ GB VRAM consumption. For the control branch, we conduct fine-tuning on the LAION-Aesthetics dataset [55], specifically selecting images with aesthetic scores above $6.5$. The training process spans $3$ epochs with a learning rate of $65e-$ 6 and batch size of $8$.
We choose PiDiNet [61] as the edge extractor. Fig. 10 shows that it strikes a better balance between geometric structure preservation and simulation of human-like strokes.
Figure 10: Extracted edge between different methods [61, 70, 60].
6.2 Painting Assistor
We fine-tune a LLaVA-1.5 model with $7B$ parameters for Draw&Guess task on our own constructed dataset in Sec. 3.2, leveraging LoRA [23]. The LoRA rank and alpha are $64$ and $16$ respectively. The model is trained for $3$ epochs with a learning rate of $52e-$ 5 and batch size of $8$, taking $5$ hours on 3$\times$×A6000 GPUs. Under $4$-bit quantization, the model achieves real-time prompt inference within $0.3$ seconds using only $5$ GB VRAM, enabling efficient on-the-fly prompt generation with satisfactory accuracy.
6.3 Idea Collector
Cross-platform Support: We implement the Idea Collector as a modular ReactJS component library, designed for cross-platform compatibility with various generative AI frameworks, such as Gradio [1] and ComfyUI [11]. The architecture separates client-side user interactions from server-side model computations through HTTP protocols, enabling platform-independent deployment via standard HTML rendering.
Besides Gradio, MagicQuill can also be integrated into ComfyUI as a custom node, as shown in Fig. 11, with customizable widgets for parameter settings and extensible architecture for future platform integrations.
@@ 6.3 Idea Collector
Usage Scenario: To demonstrate the user-friendly workflow of MagicQuill, we present an illustrative scenario: A user wants to modify an image of a complete cake, cutting a slice out of it, as shown in Fig 2. The user begins by uploading the image through the toolbar, which provides access to a range of tools (Fig. 2-B). Using the add brush, the user outlines the slice to be cut directly on the canvas (Fig. 2-D). Meanwhile, the Draw & Guess feature introduced in Sec. 3.2 predicts that the user intends to manipulate a “cake” and suggests the relevant prompt automatically in the prompt area (Fig. 2-A). Afterward, the user switches to the subtract brush to fill in the outlined slice, visually marking the area to be removed from the cake. For additional precision, the eraser tool is available to refine the cut. Once the adjustments are made, the user generates the image by clicking the Run button (Fig. 2-F), which runs the model detailed in Sec. 3.1.
@@
7 Comparison on MagicBrush Benchmark
We evaluate our approach against existing instruction-based image editing methods using the MagicBrush [80] benchmark. This benchmark provides high-quality image pairs for both single-round and multi-round editing scenarios, aligning well with our settings. Our comprehensive evaluation against six state-of-the-art instruction-based editing baselines [8, 72, 18, 29, 5, 73] demonstrates superior performance in both quantitative metrics and qualitative results, as show in Tab. 4 and Fig. 12.
Table 4: Quantitative comparison on MagicBrush benchmark
[…]
L1 L2 CLIP-I DINO CLIP-T L1 L2 CLIP-I DINO CLIP-T
InstructP2P 0.115 0.039 0.849 0.741 0.265 0.141 0.050 0.817 0.678 0.270
OmniGEN 0.092 0.037 0.903 0.837 0.268 0.152 0.062 0.839 0.685 0.272
SeedX 0.187 0.090 0.857 0.747 0.268 0.258 0.130 0.785 0.564 0.269
DDIM+PNP 0.100 0.026 0.858 0.785 0.278 0.131 0.039 0.824 0.709 0.281
Ledits++ 0.094 0.027 0.853 0.774 0.274 0.121 0.039 0.811 0.684 0.276
InfEdit 0.122 0.034 0.849 0.770 0.283 0.155 0.050 0.815 0.698 0.288
Ours 0.033 0.011 0.949 0.927 0.279 0.035 0.010 0.939 0.913 0.284
Figure 12: Visual comparison of editing result on MagicBrush
8 Editing Results under Complex Prompt
Our system allows users to refine the suggested prompts. Fig. 13 below shows our Editing Processor accurately reflects complex prompts.
Figure 13: Same image being edited under complex prompts.
9 Control Brush Area Size
Our UI supports brush size adjustment and an eraser to easily correct the control area. Fig. 14 shows our method generates realistic results for large-scale drawn content.
Figure 14: Editing results when user draws large area.
10 Failure Case
10.1 Failure Case of Editing Processor
Scribble-Prompt Trade-Off: We observe quality degradation when user-provided add brush strokes deviate from the semantic content specified in the prompt, a common occurrence among users with limited artistic skills. This creates a fundamental trade-off: strictly following the scribble structure may compromise the generation quality with respect to the text prompt. To address this issue, we propose adjusting the edge control strength.